# Gaussian Phase Noise Effects on Hybrid Precoding MIMO Systems for Sub-THz Transmission

### 25 Mar 2026 v1

@@ I-A Related Works
The PN impact in HP MIMO systems has been investigated using both Wiener PN models [Zhang, Faragallah, Nguyen] and GPN models [Corvaja, Corvaja2]. In [Zhang], the authors highlighted the PN effect in multi-user MIMO systems employing zero-forcing (ZF) precoding for downlink transmission under a CLO architecture. They derived a lower bound on the achievable rate based on the channel covariance matrices used for channel estimation during the uplink transmission. However, utilizing the ZF precoder requires the number of RF chains to match the number of users, which, from an energy consumption perspective, degrades energy efficiency (EE). In [Faragallah], the authors proposed several PN estimation techniques for PN compensation, with the aim of enhancing the performance of a SU-MIMO system operating with ILO architecture. They assumed perfect channel state information (CSI) knowledge, and HP matrices were designed following the approach presented in [Yu]. In [Corvaja], the authors demonstrated the performance degradation of HP SU-MIMO systems employing singular value decomposition (SVD) precoding in the presence of GPN. In [Corvaja2], they further examined the impact of channel estimation errors and GPN on system performance, assuming fixed analog beamforming matrices to reduce design complexity. Unfortunately, no theoretical expressions of the achievable rate or BER are derived in [Corvaja, Corvaja2]. Although GPN is considered the most suitable model for describing PN effects in sub-THz transmissions [Bicais2], relatively few papers in the literature address its impact on HP MIMO systems. To the best of our knowledge, no analytical BER expression has been derived for SVD-based HP SU-MIMO systems affected by GPN. This work therefore contributes such an analysis.
@@ 3. Performance Improvement without PN Compensation: Optimal Receiver and Modulation Scheme
By substituting $\Gamma=1$ and $M=4$ into (38), we can deduce the semi-analytical BER expression of $4$-QAM. Moreover, the PN-induced error floor expression for 4-QAM, assuming high-SNR and strong GPN regimes, is obtained by substituting $\Gamma=1$ and $M=4$ into (39) as follows
@@ V-A Energy Efficiency and Computational Complexity
Considering RF power amplifiers, which are closely linked to the transmitters’ energy consumption, the peak-to-average power ratio (PAPR) constitutes a crucial performance metric for communication systems. As demonstrated in [Bicaïs], the PAPR of the PQAM is a strictly increasing function of the number of amplitude levels $\Gamma$. This represents a drawback from an energy efficiency (EE) viewpoint.
For numerical simulations, the FC scheme was chosen to maximize the achievable SE. However, from an EE perspective, the FC scheme is not ideal when considering a high $N_{\text{RF}}$ value (still considering $N_{\text{RF}}\ll N_{\text{t}}$[Yu]. The PC scheme with or without switches was proposed as a compromise between SE and EE [Méndez-Rial, Ren, Qi, Ma]. The authors in [Yu] showed that increasing $N_{\text{RF}}$ rapidly degrades the EE in the FC architecture, since increasing the number of RF chains means increasing the number of PSs for analog precoding. As a result, the total power demand is expected to grow significantly. This is unlike the PC case, where a few number of PSs is connected to a group of RF chains and thereupon an EE improvement. However, for a small $N_{\text{RF}}$ value such as $N_{\text{s}}\leq N_{\text{RF}}\leq 2N_{\text{s}}-1$, as considered in our work, both EE and SE are better in FC than in PC. In short, the number of RF chains is decisive for choosing between FC or PC architecture for achieving a good EE.
@@
where $\left(\cdot\right)^{\dagger}$ denotes the Moore-Penrose pseudo-inverse. Nonetheless, the number of RF chains must be equal to the number of receive antennas and thus representing a drawback in an EE viewpoint if we assume a large number of receive antennas or a large number of users in multi-user MIMO systems.
@@ V-B Perspectives
The GSM showed its importance in full digital MIMO by providing an alternative for achieving high throughput despite the constraint of low modulation order in the presence of strong GPN [Saad2]. Additionally, GSM has already been investigated in HP MIMO system without PN impairments in terms of BER and EE [Li2]. A perspective to this work is to study the implementation of GSM in HP MIMO system impaired by GPN since there is no contribution proposed in the state-of-the-art to the best of our knowledge. Also, implementing MC waveforms such as OFDM and DFT-s-OFDM might be interesting to evaluate the possibility of reusing all 3GPP concepts on framing, definition of PHY signals, channel estimation and so on.
Furthermore, a new stacked intelligent metasurface (SIM)-based concept is proposed to achieve optimal EE performance. The authors in [Jiancheng] propose a promising transceiver architecture based on multi-layer SIM to realize precoding and combining in the wave domain. Basically, the digital and/or analog precoding stages are done in the wave domain to provide a very low hardware cost and therefore, a very high EE compared to exisiting MIMO technologies. However, according to the sub-THz properties, some constraints such as high-resolution ADC/DAC and the generated PN will certainly affect its performance. Thus, another perspective of this work is to study the performance of SIM-based MIMO systems with GPN. Furthermore, it could also be interesting to present a performance comparison with the HP MIMO system.

### 26 Mar 2026 v2

@@ I-A Related Works
The PN impact in HP MIMO systems has been investigated using both Wiener PN models [Zhang, Faragallah, Nguyen] and GPN models [Corvaja, Corvaja2]. In [Zhang], the authors highlighted the PN effect in multi-user MIMO systems employing zero-forcing (ZF) precoding for downlink transmission under a CLO architecture. They derived a lower bound on the achievable rate based on the channel covariance matrices used for channel estimation during the uplink transmission. However, utilizing the ZF precoder requires the number of RF chains to match the number of users, which, from an energy consumption perspective, degrades energy efficiency ($E$). In [Faragallah], the authors proposed several PN estimation techniques for PN compensation, with the aim of enhancing the performance of a SU-MIMO system operating with ILO architecture. They assumed perfect channel state information (CSI) knowledge, and HP matrices were designed following the approach presented in [Yu]. In [Corvaja], the authors demonstrated the performance degradation of HP SU-MIMO systems employing singular value decomposition (SVD) precoding in the presence of GPN. In [Corvaja2], they further examined the impact of channel estimation errors and GPN on system performance, assuming fixed analog beamforming matrices to reduce design complexity. Unfortunately, no theoretical expressions of the achievable rate or BER are derived in [Corvaja, Corvaja2]. Although GPN is considered the most suitable model for describing PN effects in sub-THz transmissions [Bicais2], relatively few papers in the literature address its impact on HP MIMO systems. To the best of our knowledge, no analytical BER expression has been derived for SVD-based HP SU-MIMO systems affected by GPN. This work therefore contributes such an analysis.
@@ 3. Performance Improvement without PN Compensation: Optimal Receiver and Modulation Scheme
By substituting $\Gamma=1$ and $M=4$ into (38), we can deduce the semi-analytical BER expression of 44-QAM. Moreover, the PN-induced error floor expression for 4-QAM, assuming high-SNR and strong GPN regimes, is obtained by substituting $\Gamma=1$ and $M=4$ into (39) as follows
@@ V-A Energy Efficiency and Computational Complexity
Considering RF power amplifiers, which are closely linked to the transmitters’ energy consumption, the peak-to-average power ratio (PAPR) constitutes a crucial performance metric for communication systems. As demonstrated in [Bicaïs], the PAPR of the PQAM is a strictly increasing function of the number of amplitude levels $\Gamma$. This represents a drawback from an energy efficiency ($E$) viewpoint.
For numerical simulations, the FC scheme was chosen to maximize the achievable SE. However, from an $E$ perspective, the FC scheme is not ideal when considering a high $N_{\text{RF}}$ value (still considering $N_{\text{RF}}\ll N_{\text{t}}$[Yu]. The PC scheme with or without switches was proposed as a compromise between SE and $E$ [Méndez-Rial, Ren, Qi, Ma]. The authors in [Yu] showed that increasing $N_{\text{RF}}$ rapidly degrades the $E$ in the FC architecture, since increasing the number of RF chains means increasing the number of PSs for analog precoding. As a result, the total power demand is expected to grow significantly. This is unlike the PC case, where a few number of PSs is connected to a group of RF chains and thereupon an $E$ improvement. However, for a small $N_{\text{RF}}$ value such as $N_{\text{s}}\leq N_{\text{RF}}\leq 2N_{\text{s}}-1$, as considered in our work, both $E$ and SE are better in FC than in PC. In short, the number of RF chains is decisive for choosing between FC or PC architecture for achieving a good $E$.
@@
where $\left(\cdot\right)^{\dagger}$ denotes the Moore-Penrose pseudo-inverse. Nonetheless, the number of RF chains must be equal to the number of receive antennas and thus representing a drawback in an $E$ viewpoint if we assume a large number of receive antennas or a large number of users in multi-user MIMO systems.
@@ V-B Perspectives
The GSM showed its importance in full digital MIMO by providing an alternative for achieving high throughput despite the constraint of low modulation order in the presence of strong GPN [Saad2]. Additionally, GSM has already been investigated in HP MIMO system without PN impairments in terms of BER and $E$ [Li2]. A perspective to this work is to study the implementation of GSM in HP MIMO system impaired by GPN since there is no contribution proposed in the state-of-the-art to the best of our knowledge. Also, implementing MC waveforms such as OFDM and DFT-s-OFDM might be interesting to evaluate the possibility of reusing all 3GPP concepts on framing, definition of PHY signals, channel estimation and so on.
Furthermore, a new stacked intelligent metasurface (SIM)-based concept is proposed to achieve optimal $E$ performance. The authors in [Jiancheng] propose a promising transceiver architecture based on multi-layer SIM to realize precoding and combining in the wave domain. Basically, the digital and/or analog precoding stages are done in the wave domain to provide a very low hardware cost and therefore, a very high $E$ compared to exisiting MIMO technologies. However, according to the sub-THz properties, some constraints such as high-resolution ADC/DAC and the generated PN will certainly affect its performance. Thus, another perspective of this work is to study the performance of SIM-based MIMO systems with GPN. Furthermore, it could also be interesting to present a performance comparison with the HP MIMO system.